UT 23 Aug 2020 - Night #2 of Restart

OSA: D.G. Huerta

ISp: E. Solheid

Mtn Manager: R. Hansen

ISA: B. Rothberg (Tucson - secret bunker)

Special Guest star: O. Kuhn (Tucson)


  • LBC-Blue vertical line noise seen on all chips in test images (not biases) - similar to IT 6549 - LBC-B is not released for on-sky use yet
    • Controller 4 is science, Controller 6 is Tech
    • New IT is 8161 - markes as Critical

Night Log:

UT 01:01 - Rain reported on the summit. Leaving MODS in the beam, but with M3 on SX. If we are closed we will continue to run calibrations.

UT 01:41 - LBC Housekeeping off, powering back on, powering up LBCs will take test biases

UT 01:44 - R. Hansen called - major storm approaching, moving telescope to horizon. Will power off LBCs after they are up

Red Rotator failed

2020/08/23 01:46:05.755034 E R rotator system encountered a problem: problems with profibus encoder reading (8) [src/channel/channel.c:537]
2020/08/23 01:46:05.755092 E R RotatorInitialize failed, disabling ROTATOR system
Powering off LBCs until later.

UT 03:31 - bright illuminations of roof camera - presumably these are lightning flashes

LBC closed Dome Tests

UT 03:44 - running up LBCs

UT 03:48 -running checkout Biases scripts

-looks OK so far, running 10 more on each side

UT 04:04 - RPC error on blue

2020/08/23 04:02:10.821221 OB execution error: one or more systems failed [src/lbc/actions.c:1162]
2 2020/08/23 04:02:10.770259 >>> preset/expose/save interrupted at preset time, retcode=7
1 2020/08/23 04:02:10.770153 camera system encountered a problem 6 [src/channel/channel.c:1138]
0 2020/08/23 04:02:10.770084 CAMERA RPC Client/Host communication failed (4) [src/camera/camera.c:1169]

Hit play again, same error, powering off then powering on again to try - see IT 2397

Error on powering off:

2020/08/23 04:09:56.584117 camera system encountered a problem: RPC Client/Host communication failed (6) [src/channel/channel.c:702]

UT 04:11 - powering LBCs back on again

UT 04:14 - Blue came back on happy

UT 04:15 - running 10 Bias checkout again and see if insanity ensues...

UT 04:20 - noticing some noise on red biases (bottom of each chip)
UT 04:22- set of 10 ran fine, running another set

- telescope is back at zenith

UT 04:29 - running Allfilters_PAO script - cycles through all flters, we have lower floods on.

Tracker images /images/tftp on CMU

lbcb.2200823.042953.fits - first blue image - electonic noise streaks on the left side of each chip

-stoppting script after 4exposures for each LBC - Red looks ok

1 for LBC-B 3 for LBC-R

-made a mono LBC-B script - 3x45 sec U-bessel and 3x10 sec g-sloan to test to see if same detector oddity shows up

UT 04:48 - executing test script

-see similar - more pronounced on chip4 - lbcb.20200823.044852.fits; lbcb.20200823.045018; lbcb.20200823.045143.fits - few thousand counts

lbcb.20200823.045308.fits - g-sloan 40k counts - shows similar streaks in all chips as seen in the first image lbcb.220 0823.042953.fits (Uspec)

lbcb.20200823.045357.fits - same as above

lbcb.20200823.045445.fits - same as above

UT 04:58 - powering off LBCs, then will power back on

V2 of script is 1sec for g-sloan and 45sec for U-Bessell

IT 6549 looks to be same problem - bad controller

-images show same thing again, up to 8400 counts, shows the most striping on Chip4 (lbcb.20200823.050612.fits) -45sec U-Bessel

-lbcb.20200823.051107 - g-sloan - 1sec - 5400 counts max - see it on all chips except 2 (chip 4 is still the worst)

UT 05:13 - powering down LBCs -

MODS Closed Dome Tests

UT 05:38 - Running up MODS-1 to test the issue with the Neon lamp. We had an issue with the MODS2 Ne lamp dangling out of the socket (IT 7869). It only intermittently came on, but there was a continuum spectrum seen. That does not appear to have been the case in this instance - when it did not work, the spectrum looked like a bias. But we'll try again and try with a configuration that might see stray light. Using MODS-2 as check.

UT 05:51 - Taking sieveSnaps. MODS1 timed out on CALMODE but a retry worked. Spots on mods1b are a bit elongated in the x direction.

UT 06:01 - Ne lamp, dual grating mode, 0.6" slit. 0002 The MODS1 Ne lamp came on. The MODS2 Ne lamp is a lot weaker and the Ne lines in the blue channel barely show up. The counts in the bluest bright line in the MODS2 spectrum are about 1200 while they are 5000 for MODS1.

UT 06:08 Ne lamp, dual grating, 0.6. 0003 MODS1 settings unchanged. MODS2 using 4x previous exposure times (blue 8-sec and red 2-sec). But now the line has 12k counts. Perhaps the lamp changes intensity with time.

UT 06:11 For MODS2, back to 2 and 0.5-sec. 0004 The line at X= 6188 has 1500 counts.

UT 06:17 MODS2 Red, 1-sec, m2r 0005 3000 counts.

UT 06:21 MODS2 Red, 2-sec again m2r 0006 once again, the line now has about 12k counts. The counts have quadrupled but the exposure time only doubled. ?? This line is at the blue end of red, perhaps in the dichroic cross-over region, but I am not sure how that could cause the scaling we see. Will look at other lines, nevertheless, and I think we may want to increase the exposure time for the MODS2 Ne (red-only) from 0.5 to 1sec. MODS2 Ne+Hg (dual) LS lamps should probably remain at 0.5, though at 1-sec it is not bad, only saturating 2 lines.

MODS focus sequences:

Time Mode M1B M1R M2B M2R
UT 06:47 Dual Imaging 7-17 7-17 7-17 9-19
UT 07:14 Dual Grating 18-28 18-28 18-28 20-30
UT 07:49 Red Grating   29-39   31-41
UT 08:58 Red Imaging   40-53   42-55
UT 09:20 Blue Grating 29-39   29-39  
UT 09:55 Dual Prism 40-50 54-64 40-50 56-66
UT 10:19 Blue Imaging 51-61   51-61  
UT 10:45 Red Prism   65-75   67-77
UT 11:04 Blue Prism 62-72   62-72  
UT 10:22 - MODS2 blue calmode error timed out after 90sec, resent again, OK

UT 11:28 - Finished with focus scripts - putting both MODS to sleep
Enclosure Open

UT 08:38 - opening enclosure for MODS-Bino Check

-will go to StoneO first for AGW Rotator ZP Check

UT 08:45 - right shutter door of enclosure is not opening

-DG.Huerta going up to check

UT 08:55 - issues appears resolved, doors closed, band of couds coming in - can resume calibrations (so close!)


  • Telescope Wake-up ~0.5hr
    • test presets (sidereal and NS)
  • Test SX AdSec out of the beam with LBC-Blue in the beam
  • LBC Wake-up -
    • Collimation - DoHybrid, then DOFPIA (~1hr)
    • Superfocus Sequence, PA=0, 180 (1hr)
    • Rotator Center Checks (pairs of 0,180; 90, 270) ~0.5hr
    • Test Science OB ~ 1hr
      • Sidereal
      • Non-Sidereal
    • Photometric Fields and Spectrophotometric Stars (to establish ZPs) ~ 1-1.5hrs
      • Requires multiple airmass visits (clear conditions, any seeing)
      • Cycle through all filters
      • This is now a high priority because we need to define ZPs
    • Twilight Flats
  • MODS
    • Basic Checkout
      • Verify collimation
      • Verify AGW 5/6 Hotspots are centered
      • Rotator ZP verification
      • AGW Transform if deemed necessary (i.e. Rot ZPs significantly changed or dithering offsets found to be problematic)
      • Observe a standard star in all modes (30-45 min for 3 images). If clear, obtain the full set of 3 images per mode; if not, make a note in the object name and obtain only 1 image per mode.
      • Dice-5 imaging offsets
    • Test Science OB
      • Sidereal
      • Non-Sidereal

Closed Dome Activities

  • MODS Prism Calibrations
    • Dual
    • Red
    • Blue
  • LBC Biases
  • Other?
-- BarryRothberg - 22 Aug 2020
Topic revision: r10 - 23 Aug 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback